Skip to end of metadata
Go to start of metadata

Regular operations today with soundings at 0, 6, 12, 18Z.  Weather warm and light variable winds, thin upper level cloud.

The NCAS profiler is back up (see separate blog entry) after the trailer computer stalled.  The sounding DM computer rebooted this morning for unknown reasons, although the UPS was heard to beep briefly during the noon sounding so perhaps there were power fluctuations. 

Briefly paused 1290 Wind Profiler at 17UT to adjust RASS range since it's getting hot.  Shifted 5C warmer, so spect sig now 5C-35C, source 0C-40C.

  • No labels

1 Comment

  1. William Brown AUTHOR

    Gary says the time stamp on the DM skipped briefly backwards for unknown reasons, perhaps ripples in the space-time continuum 8-)